Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#230 closed flight processing (fixed)

GB08/10, flight day 075a/2009, Dennys Wood

Reported by: benj Owned by: knpa
Priority: alpha 4 high Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: ~arsf/arsf_data/2009/flight_data/uk/GB08_10-2008_075a_Dennys_Wood

Data arrived from ARSF via SATA J on 19th March 2009.

Scientific details: see /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-10 Science Case.pdf

Priority: a4H

PI: Ross Hill

Sensors:

  • Eagle
  • Hawk
  • ALS-II Lidar

Quicklook

TBD.

Change History (30)

comment:1 Changed 10 years ago by benj

  • Summary changed from GB08/10, flight day 075a, Dennys Wood to GB08/10, flight day 075a/2009, Dennys Wood

comment:2 Changed 10 years ago by benj

  • Description modified (diff)
  • Priority changed from medium to high

comment:3 Changed 10 years ago by jatt

  • Owner set to jatt

comment:4 Changed 10 years ago by mark1

  • Owner changed from jatt to mark1

comment:5 Changed 10 years ago by mark1

Applanix processed ok.

Due to pre-season boresight problems with the Eagle/Hawk instruments this has been used as a 'test' set to try scripted tiepointing and boresighting per line. Appears to have been successful and results are, on the whole, of similar quality to the usual methods.

comment:6 Changed 10 years ago by mark1

Timing errors on Eagle lines:
1:GPTOFF=-0.04
3:GPTOFF=-0.07
4:GPTOFF=-0.05 SCTOFF=-0.04
5:GPTOFF=-0.05
6:GPTOFF=-0.05
7:GPTOFF=-0.05
8:GPTOFF=-0.05
9:GPTOFF=-0.03

Timing errors on Hawk lines:
1:GPTOFF=-0.05
3:GPTOFF=-0.06 SCTOFF=0.03 no sync message
4:GPTOFF=-0.05
5:GPTOFF=-0.05
6:GPTOFF=-0.05
7:GPTOFF=-0.07 SCTOFF=-0.02
8:GPTOFF=-0.05
9:GPTOFF=-0.02

comment:7 Changed 10 years ago by mark1

  • Owner changed from mark1 to jatt

Data ready to be checked in:
~mark1/scratch_space/auto_boresight/dennywood/delivery/20090408/GB08_10

comment:8 Changed 10 years ago by jatt

all looks good, rsynced to: /users/rsg/arsf/arsf_data/in_progress/2009/flight_data/uk/GB08_10-2008_075a_Dennys_Wood/

comment:9 Changed 10 years ago by mark1

Before sending needs new data quality report adding to delivery

comment:10 Changed 10 years ago by mark1

up to date data quality report added

comment:11 Changed 10 years ago by jatt

  • Owner changed from jatt to mark1

Eagle and Hawk delivered 27/04/2009

comment:12 Changed 10 years ago by mggr

  • Priority changed from alpha 5 to alpha 4 high

Corrected grading following fixing trac's priority list..

comment:13 Changed 10 years ago by mark1

corrected directory name from GB08_10-2008_075a_Dennys_Wood/ to GB08_10-2009_075a_Dennys_Wood/

comment:14 Changed 10 years ago by mark1

  • Owner mark1 deleted

comment:15 Changed 10 years ago by mark1

Deleted rcd directory from ~arsf since photos related to different project (on a different day).

comment:16 Changed 9 years ago by emca

  • Owner set to emca
  • Status changed from new to accepted

Starting LIDAR processing

comment:17 Changed 9 years ago by emca

Leica directory under ~arsf contained data relating to GB08_08 068. Removed this and replaced with the correct data for this project (copied from GB08_08 075b)

comment:18 Changed 9 years ago by emca

LIDAR processing complete and ready for delivery check at:
~airborne/workspace/GB08_10-2009_075a_Dennys_Wood/delivery/

comment:19 Changed 9 years ago by mark1

LIDAR delivery checked, data loaded and viewed in lag and looks good, everything else ok. Ready to deliver.

comment:20 Changed 9 years ago by emca

LIDAR data sent to PI 22/03/2010

comment:21 Changed 9 years ago by emca

  • Component changed from Processing: general to Archiving

LiDAR data rsynced to
/users/rsg/arsf/arsf_data/2009/flight_data/uk/GB08_10-2009_075a_Dennys_Wood

All data now delivered. Ready for archiving.

comment:22 Changed 9 years ago by knpa

  • Owner changed from emca to knpa
  • Status changed from accepted to assigned

comment:23 Changed 9 years ago by knpa

Starting archiving.

comment:24 Changed 9 years ago by knpa

No webcam data for this project.

comment:25 Changed 9 years ago by knpa

Tarball created, awaiting download by NEODC.

comment:26 Changed 9 years ago by knpa

NEODC have downloaded and backed up.

Moved to archive disk.

comment:27 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:28 Changed 6 years ago by stgo

SCTs from config file.

FlightlineEagle SCTHawk SCT
1.-0.055-0.09
2.-0.055-0.055
3.-0.055-0.025
4.-0.095-0.055
5.-0.055-0.055
6.-0.055-0.055
7.-0.055-0.075
8.-0.055-0.055
9.-0.055-0.055

comment:29 Changed 3 years ago by asm

Archiving

Everything has been archived with exception of the reprocessed hyperspectral data. Will start with this now.

comment:30 Changed 2 years ago by asm

Archiving

Reprocessed hyperspectral has been archived and it is available from CEDA at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/GB08_10/GB08_10-2009_075a_Dennys_Wood/hyperspectral-reprocessed

Plymouth 11th May 2017

Last edited 2 years ago by asm (previous) (diff)
Note: See TracTickets for help on using tickets.